OpenStudy (anonymous):

I can't understand how to set up the x/f(x) table with this information... Help Please? An animal either dies ( D ) or survives ( S ) in the course of a surgical experiment. The experiment is to be performed first with two animals. If both survive, no further trials are to be made. If exactly one animal survives, one more animal is to undergo the experiment. Finally, if both animals die, two additional animals are to be tried. (a) List the sample space to answer the next questions. (b) Assume that the trials are independent and the probability of survival in each trial is 1/4

OpenStudy (anonymous):

Assign probabilities to the elementary outcomes to answer the next question. (c) Let X denote the number of survivors. Obtain the probability distribution of X by referring to part (b). Write the x -values in ascending order. Give exact answers (in form of fraction if necessarry).

OpenStudy (anonymous):

although i am not a big fan of tree diagrams, it might be helpful in this case

OpenStudy (anonymous):

or maybe not the probability that both survive is \(\frac{1}{4}\times \frac{1}{4}=\frac{1}{16}\)

OpenStudy (anonymous):

Do you always multiply the probability like that

OpenStudy (anonymous):

good question! you do only when the trials are "independent" which you are told they are

OpenStudy (anonymous):

okay because I am given a table of 6 empty slots. 3 for x and 3 for f(x) and it says they equal 1. So the x values would be 1/16 and the f(x) one would be divided evenly into three?

OpenStudy (anonymous):

i have no idea what the table looks like, but lets see what the possibilities for the number of animals surviving are

OpenStudy (anonymous):

both could survive the first experiment, so if we say \(x\) is the number of survivors, one possibility is that \(x=2\)

OpenStudy (anonymous):

thats what the table looks like

OpenStudy (anonymous):

oooh ok i see lets go slow

OpenStudy (anonymous):

okay. so how would we start it

OpenStudy (anonymous):

first off, note that is says "Let X denote the number of survivors"

OpenStudy (anonymous):

so \(x\) is not a probability at all, it is a whole number, the number of survivors

OpenStudy (anonymous):

how many survivors can there be? in other words, what possible value can \(x\) take?

OpenStudy (anonymous):

couldn't it take any? as long as the value of f(x) equals 1

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

\(x\) denotes the number of survivors of the experiment the way the set up for the experiment is, there could be 2 survivors 1 survivor or 0 survivors so the only three possible values of \(x\) are \(2,1,0\)

OpenStudy (anonymous):

OHHH!

OpenStudy (anonymous):

that is why there are three slots on the left

OpenStudy (anonymous):

left column i mean

OpenStudy (anonymous):

got it so then it says each is independent

OpenStudy (anonymous):

now we have to figure out the probability of each value of \(x\)

OpenStudy (anonymous):

meaning the probability is 1/4 so do I multiply that by something? I sort of remember it from class but I'm not too sure

OpenStudy (anonymous):

we need 1) the probability that exactly 2 survive 2) the probability that exactly 1 survives and 3) the probability than none survive

OpenStudy (anonymous):

it is going to take a while, because there are cases to consider

OpenStudy (anonymous):

what do you mean? Aren't I supposed to set them p and 2/4 1/4 0/4

OpenStudy (anonymous):

heck no

OpenStudy (anonymous):

Shoot. Thats what my example said.

OpenStudy (anonymous):

the first column should read \(x=2,x=1,x=0\) or put 2, 1, 0 in the first column

OpenStudy (anonymous):

thats done

OpenStudy (anonymous):

now we have to consider all the ways you can end up with two live animals there are a few

OpenStudy (anonymous):

first off, both could survive the first experiment the probability of that is as computed \(\frac{1}{16}\) but there are other ways two can survive

OpenStudy (anonymous):

If exactly one animal survives, one more animal is to undergo the experiment

OpenStudy (anonymous):

so if one survives and then another one survives, you have to surviving animals the probability that exactly one survives is \[2\times \frac{1}{4}\times \frac{3}{4}=\frac{3}{8}\]

OpenStudy (anonymous):

where did you get the 1/4 and 3/4

OpenStudy (anonymous):

survival in one trial is \(\frac{1}{4}\) so not surviving is \(\frac{3}{4}\)

OpenStudy (anonymous):

okay so then for the 1 survivor. It would be 1 x 3/4 because there is only one? then for 0 the same

OpenStudy (anonymous):

you have two animals, say A and B the probability that A survives and B does not is \(\frac{1}{4}\times \frac{3}{4}\) and the probability that B survives and A does not is the same, so the probability that exactly one survives in the first round is \(2\times \frac{1}{4}\times \frac{3}{4}=\frac{3}{8}\)

OpenStudy (anonymous):

then you do the experiment again, so given that one survived, the probability that another survives, for a total of two surviving, is \(\frac{1}{8}\times \frac{3}{4}=\frac{3}{32}\)

OpenStudy (anonymous):

and finally there is yet another way to get two survivors both die on the first trial, with probability \(\frac{3}{4}\times \frac{3}{4}\) and then the next two live, with probability \(\frac{1}{4}\times \frac{1}{4}=\frac{1}{16}\)

OpenStudy (anonymous):

multiply those together, and get \[\frac{9}{256}\]

OpenStudy (anonymous):

now the total ways to get two survivors is the sum of the previous answer namely \[\frac{9}{256}+\frac{3}{32}+\frac{1}{16}=\frac{49}{256}\]

OpenStudy (anonymous):

now we have to figure out the way for exactly one to survive

OpenStudy (anonymous):

one way is that one survives the first round, and then in the second round the animal dies another way is that both die on the first round, and then one survives on the second

OpenStudy (anonymous):

any ideas how to do that?

OpenStudy (anonymous):

honestly this feels a lot more complicated than i thought. When we did it in class we would just make them equal the one. I see how you got the previous answers now to find that they both die on the first round it would be similar to the one before. Where you multiplied them.

OpenStudy (anonymous):

we already know the probability that one survives the first round is \(\frac{3}{8}\) then the second one not surviving is \(\frac{3}{4}\) so the probability that one survives the first round, and then the second one dies is \(\frac{3}{8}\times \frac{3}{4}=\frac{9}{32}\)

OpenStudy (anonymous):

probability that none survive the first round and exactly one survives the second round is \[\frac{3}{4}\times \frac{3}{4}\times \frac{3}{8}=\frac{27}{128}\]

OpenStudy (anonymous):

then that would be the answer? don't they all need the same denominator

OpenStudy (anonymous):

add those up and get \[\frac{9}{32}+\frac{27}{128}=\frac{63}{128}\]

OpenStudy (anonymous):

and finally, what is the probability that none survives?

OpenStudy (anonymous):

both die on first trial, and both die on second trial

OpenStudy (anonymous):

\[\left(\frac{3}{4}\right)^4=\frac{81}{256}\]

OpenStudy (anonymous):

Okay I think I get it I just have to keep finding the probabilities. Just like you did with the first ones.

OpenStudy (anonymous):

btw it says to write in ASCENDING order

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Lol thanks! I read that right now too

OpenStudy (anonymous):

so you should probably write 0,1,2 for the first column

OpenStudy (anonymous):

then i think the second column \[P(x=0)=f(0)=\frac{81}{256}\] then \[P(x=1)=f(1)=\frac{63}{128}\] and finally \[P(x=2)=f(2)=\frac{49}{256}\]

OpenStudy (anonymous):

lets see if they add up to 1, or if there is a mistake

OpenStudy (anonymous):

they do!

OpenStudy (anonymous):

if you multiple the middle one by 2

OpenStudy (anonymous):

looks good to me!

OpenStudy (anonymous):

oh i see what you mean, in order to add you multiply the middle one top and bottom by 2 to make the denominator 256 hope this was more or less clear it was a bit long and complicated

OpenStudy (anonymous):

Thanks a bunch. I was but I read back through your messages and it made sense! Thanks @satellite73 (:

OpenStudy (anonymous):

yw